# Dexterous Manipulation through Imitation Learning: A Survey

### v4

@@ V-A2 Mocap Gloves
Motion capture systems typically utilize stable hardware devices such as multi-camera setups with markers, IMU sensors, and RGB-D cameras. These devices are robust against changes in lighting, occlusion, and complex backgrounds. Mocap gloves collect human hand motion data directly via sensors, enabling ideal real-time performance and significantly improving data collection efficiency in teleoperation. Although motion capture gloves are expensive, they provide precise hand tracking [282]. Wang et al. [16] introduced DexCap, a portable motion capture system. It includes a mocap glove for accurate finger joint tracking, a single-view camera for 6-DoF wrist pose tracking, and an RGB-D LiDAR camera for observing the surrounding 3D environment. With this precise 3D hand motion data, the proposed DexIL system can effectively learn bimanual dexterous manipulation skills. The remote system features two Franka Emika robotic arms, each with a LEAP dexterous robotic hand. Similarly, Mosbach et al. [283] used the SenseGlove DK1, a force-feedback glove, to capture hand joint movements, with hand tracking facilitated by a camera mounted on a headset.
@@ V-A3 VR/AR Controllers
VR devices typically include a head-mounted display, a tracking system, and input devices. The head-mounted display provides an immersive visual experience with high-resolution screens and head motion tracking. The tracking system captures the user’s movements to ensure that interactions in the virtual environment correspond to real-world actions. Input devices, such as controllers or gloves, facilitate user interaction within the virtual space. Zhang et al. [14] developed a teleoperation system using consumer-grade VR devices to control a PR2 robot. Following this, methods utilizing low-cost equipment [284, 285] demonstrated high-quality teleoperation through mixed reality. To streamline scene construction, Mosbach et al. [283] explored VR teleoperation in simulated environments for manipulation tasks. Recently, Bunny-VisionPro [286] equipped the Apple VisionPro with a haptic module to provide tactile feedback. Similarly, Open-television [287] used an active camera mounted on a humanoid robot to capture first-person stereo videos. This approach enhances the robot’s ability to perform precise and context-aware actions by providing a dynamic, real-time visual perspective similar to human vision. Lin et al. [288] introduced a low-cost teleoperation system, HATO, combining two Psionic Ability Hands for prosthetic use with UR5e robot arms. The system utilizes two Meta Quest 2 VR controllers with IMU sensors to capture hand spatial positions and orientations, translating controller inputs into multi-fingered hand poses.
@@ V-A4 Exoskeleton and Bilateral Systems
Exoskeletons are wearable devices that gather and analyze user motion data. Fabian [289] developed a lightweight exoskeleton, DE VITO, for measuring human arm movements to teleoperate the mobile robot DE NIRO[290]. Meanwhile, AirExo [291] presents a framework for whole-arm dexterous manipulation adaptable to different robot arms using interchangeable 3D-printed components for robots divergent in morphology.
Another approach involves a bilateral framework, where the movements of the leader robot are mirrored by the follower robot. Any resistance or force encountered by the follower is communicated back to the leader, enabling precision and tactile sensation tasks. Kim et al. [292] develop a controller with Denavit-Hartenberg (DH) parameters matching the teleoperated dual-arm robot alongside a calibration method to reduce gravitational errors. For demonstrations without real robots, the controller uses force/torque (F/T) sensors identical to those in real robots to provide force feedback. Recently, employing cost-effective arms with comparable size as leaders and followers, ALOHA [77] utilizes structurally analogous robotic arms with identical joint spaces for teleoperation. Expanding upon this concept, Mobile ALOHA [143] integrates the system with an Automated Guided Vehicle (AGV) to establish a whole-body teleoperation system. GELLO [293] reduces costs by replacing the real robotic arms at the local site with scaled kinematically equivalent 3D-printed parts, achieving one-to-one joint mapping. After that, UMI [144] further eliminates the need for physical robot arms by using hand-held grippers, providing a more portable interface for in-the-wild data collection.
Differing from specific robot methods, AnyTeleop [19] introduced a unified system supporting multiple robot arms and dexterous hands through a general human-robot hand retargeting method. This system supports different arms by generating trajectories based on the estimated Cartesian end-effector pose. ACE [280] developed a cross-platform visual-exoskeleton teleoperation system compatible with diverse robot hardware, including various end-effectors such as grippers and multi-finger hands, offering flexibility. Its exoskeleton arm features high-resolution encoders for precise joint position readings, ensuring accurate end-effector tracking.
@@
Motion-centric Imitation Learning DexMV [309] Robotic Telekinesis [17] Track2Act [310] Unlabeled third-person real-world videos Unlabeled third-person unpaired videos Unlabeled videos, limited real-world data Reconstructs 3D hand-object trajectories or end-effector trajectories from videos and maps them to robot actions, enabling cross-domain imitation and improved generalization.
Synthetic Video for Policy Learning Gen2Act [311] NIL [312]
Text-generated unlabeled synthetic videos Diffusion-generated unlabeled visual videos
Uses language-conditioned video generation to synthesize training demonstrations, removing the need for expert data. Learns policies via perceptual similarity in generated videos, supporting large-scale training and generalization.
Representation Learning for Generalization Ag2Manip [313] Unlabeled human action videos Learns agent-agnostic action embeddings to support imitation across different robot embodiments and improve generalization.
Task-specific Architectures and Learning Objectives Bi-KVIL [314] Rank2Reward [315] ViViDex [316] Unlabeled third-person dual-hand videos Unlabeled videos Unlabeled noisy videos with no ground-truth state Specialized frameworks modeling bimanual coordination, reward learning without labels, and hierarchical learning architectures for robust imitation.
@@
Human-teleoperated Datasets MIME [138] RH20T [139] BridgeData [140] BridgeData V2 [274] 8,260 demos, 20 tasks 110K+ multimodal sequences 7,200 demos, 71 tasks, 10 domains 60,096 trajectories, 24 environments RGB-D, Kinesthetic trajectories RGB-D, Audio, Tactile, Proprioception Egocentric vision, Action sequences Vision, Language embeddings
Augmented Datasets RoboAgent [317] 7,500 → 98K* augmented trajectories RGB, Semantic masks
Synthetic Datasets MimicGen [318] 200 → 50K synthesized demos, 18 tasks Object poses, Motion primitives
Dexterous and Bimanual Manipulation Datasets ARCTIC [319] DexGraspNet [320] OAKINK2 [321] 2.1M frames (3D meshes) 1.32M grasps, 5,355 objects 627 sequences, 4.01M frames Multi-view RGB, Contact dynamics Tactile sensing, Physics simulation Multi-view capture, 3D pose annotations
@@ V-B Learning from Video Demonstrations for Dexterous Manipulation
Video-based demonstration learning offers a promising solution to the challenge of acquiring robotic manipulation skills in the absence of expensive expert demonstrations. A summary of representative approaches within this paradigm is presented in Tab. VII. By extracting structured behavioral cues from a wide range of human videos—including recorded, synthetic, and web-sourced data—these methods transform raw visual observations into robot-interpretable representations, thereby facilitating robust generalization across diverse tasks, objects, and embodiments.
@@ V-B1 Motion-centric Imitation Learning
A major line of work adopts motion-centric imitation, where human hand–object interactions are converted into robot-executable trajectories. Qin et al. [309] propose a method that reconstructs 3D hand–object poses from third-person human videos and re-targets them to dexterous robot hardware, enabling high-fidelity policy learning without the need for direct teleoperation or egocentric demonstrations. Building upon this approach, Bharadhwaj et al. [310] predict object-relative end-effector trajectories from unstructured videos and fine-tune residual policies on robots, thereby achieving generalization across novel scenes and object configurations with minimal real-world training. Complementary to these approaches, Sivakumar et al. [17] directly map unstructured human hand motions from third-person videos to dexterous robotic hands, supporting real-time teleoperation without paired demonstrations.
@@ V-B2 Synthetic Video-driven Imitation Learning
Another direction leverages synthetic video as training signals for policy learning. Bharadhwaj et al. [311] propose a method that generates realistic manipulation videos from textual prompts and internet instructions, which are then employed to train robot policies capable of generalizing to unseen categories and actions. Building upon this approach, Albaba et al. [312] introduce a diffusion-based video generation framework that replaces explicit action labels with perceptual similarity, providing a scalable pathway for training robots with diverse morphologies. By eliminating the reliance on real or expert demonstrations, these approaches significantly expand the scalability of video-based demonstration learning.
@@ V-B3 Representation Learning for Generalization
A third stream of research focuses on representation learning for generalization, aiming to decouple policy learning from specific tasks and embodiments. Singh et al. [322] propose a method that spatially aligns human hands and objects in 3D, producing consistent motion patterns that can pretrain general-purpose manipulation policies and enable task-agnostic skill acquisition. Building upon this idea, Li et al. [313] learn shared agent-agnostic visual–action embeddings from diverse human operation videos, achieving robust sim-to-real transfer even on previously unseen tasks and object configurations. These approaches highlight the potential of abstracting beyond raw trajectory reconstruction to acquire transferable skills across domains.
@@
Annotation Quality Accuracy and consistency of labels across modalities and tasks. Object pose precision, end-effector position, task boundary clarity
@@ V-B4 Task-specific Architectures and Objectives
Finally, several task-specific architectures and learning objectives have been designed to address unique challenges in video-based imitation. Gao et al. [314] introduce a hybrid master–slave modeling strategy with low-level geometric constraints for dual-hand coordination, enabling robust reproduction of bimanual manipulation in cluttered environments. Building upon this idea, Shaw et al. [323] extract vision–action–force relations from human demonstrations to provide physical and semantic priors for dexterous robot learning, reducing the dependence on paired demonstrations. Extending this approach, Chen et al. [316] present a three-stage framework that combines video demonstrations, reinforcement learning-based trajectory optimization, and vision-only policy learning without privileged object information, thus enabling physically executable policies from noisy demonstrations. Yang et al. [315] address the reward specification bottleneck by constructing progress-based rewards from temporal rankings of video frames, supporting both reinforcement and adversarial imitation without access to action annotations. Collectively, these approaches advance the practicality of video-based demonstration learning by tackling challenges such as bimanual coordination, noisy inputs, reward learning, and task-specific constraints.
@@ V-C1 Human-teleoperated Datasets
A straightforward approach to collecting data for IL is through human teleoperation, where operators directly control robots to perform tasks. MIME [138] represents an early large-scale effort in this direction, providing 8,260 human–robot demonstrations across 20 diverse tasks, along with both human demonstration videos and kinesthetic robot trajectories. RH20T [139] extends this paradigm by scaling up to over 110,000 multimodal manipulation sequences and incorporating richer sensing modalities—visual, tactile, audio, and proprioceptive—captured through teleoperation interfaces with force–torque sensors and haptic feedback. This multimodal design enables one-shot IL across a wider range of tasks, robots, and environments. Building on the same teleoperation foundation, BridgeData [140] focuses on cross-environment generalization, offering 7,200 demonstrations across 71 tasks in 10 environments, primarily in kitchen settings. Its successor, BridgeData V2 [274] expands the dataset to 60,096 trajectories in 24 environments, covering tasks from pick-and-place to complex manipulations, and supporting scalable robot learning with multi-task and language-conditioned objectives.
@@ V-C2 Augmented Datasets
To overcome the time-consuming and labor-intensive process of collecting large numbers of human demonstrations for IL, some datasets leverage data augmentation to expand existing demonstrations. RoboAgent [317] begins with 7,500 teleoperated trajectories and scales them up to roughly 98,000 using semantic augmentations, eliminating the need for additional human or robot effort. CyberDemo [324] adopts a similar strategy, collecting demonstrations in both simulated and real environments before applying extensive augmentations that introduce visual and physical variations, thereby improving policy robustness and generalization.
@@ V-C3 Synthetic Datasets
Some datasets employ simulated or synthetic demonstration generation systems to greatly expand the amount of training data from a small set of human demonstrations. MimicGen [318] generates over 50,000 demonstrations across 18 tasks from roughly 200 human examples by adapting object-centric manipulation behaviors to new contexts via trajectory transformations. IntervenGen [325] builds on this idea by autonomously producing large sets of corrective interventions from minimal human input, thereby improving policy robustness to distribution shifts. DiffGen [326] adopts a similar strategy but incorporates differentiable physics simulation, photorealistic rendering, and vision–language models to create realistic robot demonstrations directly from text instructions.

### v5

@@ V-A2 Mocap Gloves
Motion capture systems typically utilize stable hardware devices such as multi-camera setups with markers, IMU sensors, and RGB-D cameras. These devices are robust against changes in lighting, occlusion, and complex backgrounds. Mocap gloves collect human hand motion data directly via sensors, enabling ideal real-time performance and significantly improving data collection efficiency in teleoperation. Although motion capture gloves are expensive, they provide precise hand tracking [caeiro2021systematic]. Wang et al. [wang2024dex] introduced DexCap, a portable motion capture system. It includes a mocap glove for accurate finger joint tracking, a single-view camera for 6-DoF wrist pose tracking, and an RGB-D LiDAR camera for observing the surrounding 3D environment. With this precise 3D hand motion data, the proposed DexIL system can effectively learn bimanual dexterous manipulation skills. The remote system features two Franka Emika robotic arms, each with a LEAP dexterous robotic hand. Similarly, Mosbach et al. [mosbach2022accelerating] used the SenseGlove DK1, a force-feedback glove, to capture hand joint movements, with hand tracking facilitated by a camera mounted on a headset.
@@ V-A3 VR/AR Controllers
VR devices typically include a head-mounted display, a tracking system, and input devices. The head-mounted display provides an immersive visual experience with high-resolution screens and head motion tracking. The tracking system captures the user’s movements to ensure that interactions in the virtual environment correspond to real-world actions. Input devices, such as controllers or gloves, facilitate user interaction within the virtual space. Zhang et al. [zhang2018deep] developed a teleoperation system using consumer-grade VR devices to control a PR2 robot. Following this, methods utilizing low-cost equipment [arunachalam2023holo, radosavovic2023real] demonstrated high-quality teleoperation through mixed reality. To streamline scene construction, Mosbach et al. [mosbach2022accelerating] explored VR teleoperation in simulated environments for manipulation tasks. Recently, Bunny-VisionPro [ding2024bunny] equipped the Apple VisionPro with a haptic module to provide tactile feedback. Similarly, Open-television [cheng2024opentelevision] used an active camera mounted on a humanoid robot to capture first-person stereo videos. This approach enhances the robot’s ability to perform precise and context-aware actions by providing a dynamic, real-time visual perspective similar to human vision. Lin et al. [lin2024learning] introduced a low-cost teleoperation system, HATO, combining two Psionic Ability Hands for prosthetic use with UR5e robot arms. The system utilizes two Meta Quest 2 VR controllers with IMU sensors to capture hand spatial positions and orientations, translating controller inputs into multi-fingered hand poses.
@@ V-A4 Exoskeleton and Bilateral Systems
Exoskeletons are wearable devices that gather and analyze user motion data. Fabian [falck2019vito] developed a lightweight exoskeleton, DE VITO, for measuring human arm movements to teleoperate the mobile robot DE NIRO[falck2018human]. AirExo [fang2024airexo] presented a framework for whole-arm dexterous manipulation adaptable to different robot arms using interchangeable 3D-printed components for robots divergent in morphology.
@@
Motion-centric Imitation Learning DexMV [qin2022dexmv] Robotic Telekinesis [sivakumar2022robotic] Track2Act [bharadhwaj2024track2act] Unlabeled third-person real-world videos Unlabeled third-person unpaired videos Unlabeled videos, limited real-world data Reconstructs 3D hand-object trajectories or end-effector trajectories from videos and maps them to robot actions, enabling cross-domain imitation and improved generalization.
Synthetic Video for Policy Learning Gen2Act [bharadhwaj2024gen2act] NIL [albaba2025nil] Text-generated unlabeled synthetic videos Diffusion-generated unlabeled visual videos Uses language-conditioned video generation to synthesize training demonstrations, removing the need for expert data. Learns policies via perceptual similarity in generated videos, supporting large-scale training and generalization.
Representation Learning for Generalization Ag2Manip [li2024ag2manip] Unlabeled human action videos Learns agent-agnostic action embeddings to support imitation across different robot embodiments and improve generalization.
Task-specific Architectures and Learning Objectives Bi-KVIL [gao2024bi] Rank2Reward [yang2024rank2reward] ViViDex [chen2024vividex] Unlabeled third-person dual-hand videos Unlabeled videos Unlabeled noisy videos with no ground-truth state Specialized frameworks modeling bimanual coordination, reward learning without labels, and hierarchical learning architectures for robust imitation.
Another approach involves a bilateral framework, where the movements of the leader robot are mirrored by the follower robot. Any resistance or force encountered by the follower is communicated back to the leader, enabling precision and tactile sensation tasks. Kim et al. [kim2023training] developed a controller with Denavit-Hartenberg parameters matching the teleoperated dual-arm robot alongside a calibration method to reduce gravitational errors. For demonstrations without real robots, the controller uses force/torque sensors identical to those in real robots to provide force feedback. Recently, employing cost-effective arms with comparable size as leaders and followers, ALOHA [zhao2023learning] utilized structurally analogous robotic arms with identical joint spaces for teleoperation. Expanding upon this concept, Mobile ALOHA [fu2024mobile] integrated the system with an automated guided vehicle to establish a whole-body teleoperation system. GELLO [wu2024gello] reduced costs by replacing the real robotic arms at the local site with scaled kinematically equivalent 3D-printed parts, achieving one-to-one joint mapping. After that, UMI [chi2024universal] further eliminated the need for physical robot arms by using hand-held grippers, providing a more portable interface for in-the-wild data collection.
Differing from specific robot methods, AnyTeleop [qin2023anyteleop] introduced a unified system supporting multiple robot arms and dexterous hands through a general human-robot hand retargeting method. This system supports different arms by generating trajectories based on the estimated Cartesian end-effector pose. ACE [yang2024ace] developed a cross-platform visual-exoskeleton teleoperation system compatible with diverse robot hardware, including various end-effectors such as grippers and multi-finger hands, offering flexibility. Its exoskeleton arm features high-resolution encoders for precise joint position readings, ensuring accurate end-effector tracking.
@@
Human-teleoperated Datasets MIME [sharma_multiple_2018] RH20T [fang_rh20t_2024] BridgeData [ebert_bridge_2022] BridgeData V2 [walke2023bridgedata] 8,260 demos, 20 tasks 110K+ multimodal sequences 7,200 demos, 71 tasks, 10 domains 60,096 trajectories, 24 environments RGB-D, Kinesthetic trajectories RGB-D, Audio, Tactile, Proprioception Egocentric vision, Action sequences Vision, Language embeddings
Augmented Datasets RoboAgent [bharadhwaj2024roboagent] 7,500 → 98K* augmented trajectories RGB, Semantic masks
Synthetic Datasets MimicGen [mandlekar2023mimicgen] 200 → 50K synthesized demos, 18 tasks Object poses, Motion primitives
Dexterous and Bimanual Manipulation Datasets ARCTIC [fan2023arctic] DexGraspNet [wang2023dexgraspnet] OAKINK2 [2403.19417] 2.1M frames (3D meshes) 1.32M grasps, 5,355 objects 627 sequences, 4.01M frames Multi-view RGB, Contact dynamics Tactile sensing, Physics simulation Multi-view capture, 3D pose annotations
@@ V-B Learning from Video Demonstrations for Dexterous Manipulation
Video-based demonstration learning offers a promising solution to the challenge of acquiring robotic manipulation skills in the absence of expensive expert demonstrations. Table VII presents a summary of representative methods in this paradigm. By extracting structured behavioral cues from a range of human videos including recorded, synthetic, and web-sourced data, these methods transform raw visual observations into robot-interpretable representations, facilitating robust generalization across diverse tasks, objects, and embodiments.
@@ V-B1 Motion-centric Imitation Learning
A major line of work adopt motion-centric imitation, where human hand–object interactions were converted into robot-executable trajectories. Qin et al. [qin2022dexmv] proposed a method that reconstructed 3D hand–object poses from third-person human videos and re-targeted them to dexterous robot hardware, enabling high-fidelity policy learning without the need for direct teleoperation or egocentric demonstrations. Building upon this approach, Bharadhwaj et al. [bharadhwaj2024track2act] predicted object-relative end-effector trajectories from unstructured videos and fine-tune residual policies on robots, thereby achieving generalization across novel scenes and object configurations with minimal real-world training. Complementary to these approaches, Sivakumar et al. [sivakumar2022robotic] directly mapped unstructured human hand motions from third-person videos to dexterous robotic hands, supporting real-time teleoperation without paired demonstrations.
@@ V-B2 Synthetic Video-driven Imitation Learning
Another direction leverage synthetic video as training signals for policy learning. Bharadhwaj et al. [bharadhwaj2024gen2act] proposed a method that generated realistic manipulation videos from textual prompts and internet instructions, which were then employed to train robot policies capable of generalizing to unseen categories and actions. Building upon this approach, Albaba et al. [albaba2025nil] introduced a diffusion-based video generation framework that replaced explicit action labels with perceptual similarity, providing a scalable pathway for training robots with diverse morphologies. By eliminating the reliance on real or expert demonstrations, these approaches significantly expanded the scalability of video-based demonstration learning.
@@ V-B3 Representation Learning for Generalization
A third stream of research focuses on representation learning for generalization, aiming to decouple policy learning from specific tasks and embodiments. Singh et al. [singh2024hand] proposed a method that spatially aligned human hands and objects in 3D, producing consistent motion patterns that could pretrain general-purpose manipulation policies and enable task-agnostic skill acquisition. Building upon this idea, Li et al. [li2024ag2manip] learned shared agent-agnostic visual–action embeddings from diverse human operation videos, achieving robust sim-to-real transfer even on previously unseen tasks and object configurations. These approaches highlighted the potential of abstracting beyond raw trajectory reconstruction to acquire transferable skills across domains.
@@
Annotation Quality Accuracy, and consistency of labels across modalities and tasks. Object pose precision, end-effector position, task boundary clarity
@@ V-B4 Task-specific Architectures and Objectives
Finally, several task-specific architectures and learning objectives have been designed to address unique challenges in video-based imitation. Gao et al. [gao2024bi] introduced a hybrid master–slave modeling strategy with low-level geometric constraints for dual-hand coordination, enabling robust reproduction of bimanual manipulation in cluttered environments. Building upon this idea, Shaw et al. [shaw2023videodex] extracted vision–action–force relations from human demonstrations to provide physical and semantic priors for dexterous robot learning, reducing the dependence on paired demonstrations. Extending this approach, Chen et al. [chen2024vividex] presented a three-stage framework that combined video demonstrations, reinforcement learning-based trajectory optimization, and vision-only policy learning without privileged object information, thus enabling physically executable policies from noisy demonstrations. Yang et al. [yang2024rank2reward] addressed the reward specification bottleneck by constructing progress-based rewards from temporal rankings of video frames, supporting both reinforcement and adversarial imitation without access to action annotations. Collectively, these approaches advanced the practicality of video-based demonstration learning by tackling challenges such as bimanual coordination, noisy inputs, reward learning, and task-specific constraints.
@@ V-C1 Human-teleoperated Datasets
A straightforward approach to collecting data for IL is through human teleoperation, where operators directly control robots to perform tasks. MIME [sharma_multiple_2018] represented an early large-scale effort in this direction, providing 8,260 human–robot demonstrations across 20 diverse tasks, along with both human demonstration videos and kinesthetic robot trajectories. RH20T [fang_rh20t_2024] extended this paradigm by scaling up to over 110,000 multimodal manipulation sequences and incorporating richer sensing modalities—visual, tactile, audio, and proprioceptive—captured through teleoperation interfaces with force–torque sensors and haptic feedback. This multimodal design enabled one-shot IL across a wider range of tasks, robots, and environments. Building on the same teleoperation foundation, BridgeData [ebert_bridge_2022] focused on cross-environment generalization, offering 7,200 demonstrations across 71 tasks in 10 environments, primarily in kitchen settings. Its successor, BridgeData V2 [walke2023bridgedata] expanded the dataset to 60,096 trajectories in 24 environments, covering tasks from pick-and-place to complex manipulations, and supporting scalable robot learning with multi-task and language-conditioned objectives.
@@ V-C2 Augmented Datasets
To overcome the time-consuming and labor-intensive process of collecting many human demonstration data for IL, some datasets leverage data augmentation to expand existing demonstrations. RoboAgent [bharadhwaj2024roboagent] began with 7,500 teleoperated trajectories and scaled them up to roughly 98,000 using semantic augmentations, eliminating the need for additional human or robot effort. CyberDemo [wang2024cyberDemo] adopted a similar strategy, collecting demonstrations in both simulated and real environments before applying extensive augmentations that introduced visual and physical variations, improving policy robustness and generalization.
@@ V-C3 Synthetic Datasets
Some datasets employ simulated or synthetic demonstration generation systems to expand the amount of training data from a small set of human demonstrations. MimicGen [mandlekar2023mimicgen] generated over 50,000 demonstrations across 18 tasks from roughly 200 human examples by adapting object-centric manipulation behaviors to new contexts via trajectory transformations. IntervenGen [2405.01472] built on this idea by autonomously producing large sets of corrective interventions from minimal human input, thereby improving policy robustness to distribution shifts. DiffGen [2405.07309] adopted a similar strategy but incorporated differentiable physics simulation, photorealistic rendering, and vision–language models to create realistic robot demonstrations directly from text instructions.